1) solve the cauchy - Euler initial value problem X²y-sty tsy :o 4cl) = 1, Y (1)-9
Answer #1

Put x = Z then 2²82 D= d. *y–5xy+ 5y = 0, yu=1, y(-) = 9 (x- Sx + y) = 0 y x=8² > log x== x²0² ~ D (D-1) ado a & : (010-1)